A coal contains 78% carbon, 4% hydrogen, 2% oxygen, 1.8%


sulphur and rest as non-combustibles.


Calculate:


(i) Gross calorific value of the coal in GJ/ton using Dulong’s Formula


337 C + 1442 [H – (O/8)] + 93 S kJ/kg


where C, H, O and S are percentages of carbon, Hydrogen, oxygen and


sulphur respectively.


(ii) Amount of coal needed per day to burn in a 10 MW power plant


working with 32% thermal efficiency (Given 1 kWh = 3.6 MJ).

1
Expert's answer
2022-04-25T20:32:02-0400

(i) HCV = 337 C + 1442 [H – (O/8)] + 93 S = 337 × 0.78 + 1442 [0.04 – (0.02/8)] + 93 × 0.018 = 318.609 kJ/kg or 0.319 GJ/ton

(ii) Q` = W/"\eta" = 10 MW / 0.32 = 31.25 MW

Q = Q`∆t = (31.25 MJ/s)(24×3600 s) = 2.7×106 MJ

mcoal = Q/q = 2.7×106 MJ / 0.319 MJ/kg = 8.46×106 kg

m'coal = mcoal / ∆t = (8.46×106 kg) / (24×3600 s) = 97.96 kg/s